OpenStudy (anonymous):

tan(45+a)+tan(45-a)=(2)/(cos2a)

OpenStudy (aravindg):

Open up tan(45+1) and tan (45-a)

OpenStudy (aravindg):

*45+a

OpenStudy (anonymous):

im up to (1+tanA)/(1-tanA) + (1-tanA)/(1+tanA)

OpenStudy (anonymous):

\[\tan(x + y) = \frac{\tan(x) + \tan(y)}{1 - \tan(x)\tan(y)}\] \[\tan(x) - \tan(y) = \frac{\tan(x) - \tan(y)}{1 + \tan(x)\tan(y)}\]

OpenStudy (anonymous):

*tan(x-y) on LHS..

OpenStudy (anonymous):

so whats next?

OpenStudy (anonymous):

Take LCM first and combine the two terms by making their denominator equal..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{1 +\tan(a)}{1 - \tan(a)} + \frac{1 - \tan(a)}{1 + \tan(a)}\]

OpenStudy (anonymous):

yeah next step please

OpenStudy (anonymous):

\[\implies \frac{(1 + \tan(a))^2 + (1 - \tan(a))^2 }{(1 - \tan(a))(1 + \tan(a))}\]

OpenStudy (anonymous):

yes and after that?

OpenStudy (anonymous):

Just expand the squares and cancel out what is being cancelled..

OpenStudy (anonymous):

ok